OpenStudy (anonymous):

integral x+2 / x^2+2x+2

hartnn (hartnn):

since the derivative of denominator is 2x+2 i suggest you write numerator as (1/2) (2x+2 +2) then separate out the denominator.

hartnn (hartnn):

did u get what adjustment i did ?

hartnn (hartnn):

so first term you will get the form as int f'(x)/f(x) dx which is log |f(x)| +c and 2nd term, you'll have as 1/x^2+2x+2 which is of the form 1/(x^2+a^2) [by completing the square in denominator] and you will get the form of tan inverse function

OpenStudy (anonymous):

{(x^2+a^2) by completing the square } do this step

hartnn (hartnn):

x^2+2x+2 = x^2+2x+1 +1 = (x+1)^1 +1^2 form ---->x^2+a^2----->1/a tan inverse (x/a) (x+1)^1 +1^2 ------> tan inverse (x+1) and don't forget the 1/2 as the co-efficient :)

OpenStudy (anonymous):

x^2+2x+2 = x^2+2x+1 +1 = (x+1)^1 +1^2 please explain this step

hartnn (hartnn):

i just split 2 as 1+1 because we know that x^2+2x+1 is a perfect square.

OpenStudy (anonymous):

ohhhhhh sorry my mistake i didn't see it correctly

hartnn (hartnn):

tell me if you get stuck :)

OpenStudy (anonymous):

no thank you you are my savior

hartnn (hartnn):

you're welcome ^_^